### v1

@@ 1 Introduction
The rapid growth of wireless technologies, driven by mobile devices, Internet-of-Things (IoT) deployments, and emerging spectrum-hungry applications, has intensified pressure on traditionally static spectrum allocation policies [4, 11, 25]. To address spectrum underutilization, database-driven cognitive radio networks (DB-CRNs) and dynamic spectrum sharing have been introduced, enabling Secondary Users (SUs) to opportunistically access licensed spectrum owned by Primary Users (PUs) through regulatory-approved geolocation databases (DBs). A prominent example is the U.S. Citizens Broadband Radio Service (CBRS), which operates in the $3.5$ GHz band originally allocated to federal and satellite services [4, 24]. Moreover, dynamic spectrum sharing is expected to be a foundational element in future networks such as 6G systems [36].
While this paradigm significantly improves spectrum efficiency, it also introduces a wide range of security and privacy challenges. In particular, regulatory authorities (e.g., FCC and ITU) require users to continuously report sensitive operational information, including precise location, device identity, and transmission parameters, to centralized DBs. This requirement raises serious concerns regarding user privacy, anonymity, and long-term profiling [27, 22]. Moreover, the location-centric design of DB-CRNs makes them inherently vulnerable to location spoofing attacks, in which malicious or compromised users falsify their positions to obtain unauthorized spectrum access [47]. At the same time, the centralized reliance on spectrum databases, the broadcast nature of wireless communication, and the proliferation of low-cost devices further expose DB-CRNs to denial-of-service (DoS) attacks that can disrupt spectrum coordination and degrade network availability [14].
@@
    We develop and evaluate a novel and efficient framework, namely SLAPX, that leverages advanced cryptographic primitives to simultaneously address multiple, often conflicting, objectives in DB-CRNs. The proposed design achieves user location privacy and anonymity during spectrum querying and access while complying with enforced regulations. This paper extends our SLAP framework [^1] The preliminary version of our approach, namely SLAP, was presented and published at the IEEE SVCC 2025 Conference [15]., incorporating substantial new material and the following key enhancements and additional features. Designed for realistic adversarial settings, SLAPX remains robust against location spoofing and DoS attacks by malicious or compromised users. The key properties of SLAPX are outlined as follows: (i) Efficient Location Privacy-Preserving and Anonymous Spectrum Query: SLAPX provides strong anonymity and location privacy throughout the spectrum access workflow while remaining fully compliant with protocol and regulatory requirements. (ii) Adaptive Location Proof and Spoofing Resistance for DB-CRN: SLAPX comprises an adaptive location proof algorithm with dual scenario support and architectural flexibility that operates under realistic network assumptions with malicious or compromised users. (iii) DoS Mitigation with DB-CRN Architecture Compliance: SLAPX provides comprehensive DoS mitigation through a proactive defense mechanism that combines Verifiable Delay Functions (VDFs) with rate limiting enforced via the linkability properties of Revocable-iff-linked Linkable Ring Signatures (RLRSs).
@@
$\bullet$ Improvements over SLAP Framework: This paper extends our SLAP framework [^2] This paper is an extended version of our paper published at IEEE SVCC 2025 Conference [15], incorporating substantial new material and the following key enhancements and additional features:
@@ 2.1 System Architecture: DB-CRN
Network Components: The DB-CRN is a spectrum-sharing architecture in which unlicensed devices query trusted geolocation databases to obtain channel availability at their current location [36]. This model was deployed in real FCC TV White Space (TVWS) systems and adopted in the Microsoft and Google database trials [10, 16]. The DB-CRN architecture consists of five key entities [35]:
@@
  • Spectrum Regulator: It is the regulatory authority that defines technical rules for DB-CRN operation, like incumbent protection requirements, device reporting and registration procedures, and transmit parameters [4]. This entity authorizes and oversees geolocation database administrators and provides authoritative incumbent datasets and regulatory constraints. Throughout this manuscript, the FCC is used interchangeably to denote this spectrum regulatory authority for convenience.
@@
  • Servers: They represent external network or cloud services (e.g., CRN, cloud, edge nodes) that users access after obtaining spectrum authorization [22]. They do not participate in spectrum assignment; instead, they simply carry user traffic over the channels approved by the PSD.
@@
  • Private Spectrum Databases (PSDs): PSDs are FCC-certified, cloud-based geolocation databases (e.g., Google, Microsoft, Spectrum Bridge) responsible for maintaining incumbent information, propagation constraints, and channel availability maps [12]. PSDs process spectrum-availability queries submitted by users and return permissible channels and power limits, synchronizing regularly according to FCC mandates [24].
@@
  • User Equipments (UEs): UEs are unlicensed devices of users (e.g., smartphones, sensors, and vehicular radios) that seek opportunistic access to licensed spectrum bands [22]. Acting as SUs, they determine their location, submit channel-availability queries to PSDs, and operate strictly within the constraints specified in PSD responses to avoid interfering with PUs. Throughout this paper, the terms client and UE are used interchangeably.
@@
  • Access Points (APs): APs consist of WiFi access points or cellular base stations (LTE/5G macro or small cells) that provide the IP connectivity required for UEs to reach PSDs and servers over the Internet [15]. APs do not manage spectrum availability but offer backhaul connectivity and, in some advanced architectures, assist with localization or timing support.
@@
Communication Flow and Protocol Stack: In a DB-CRN, UEs communicate with PSDs over standard Internet paths to obtain spectrum availability, while servers and APs provide network connectivity and application-layer services. Communication exchanges follow an IP-based client–server model and use secure, lightweight protocols consistent with real FCC TVWS deployments and the IETF PAWS specification [30]. The communication workflow consists of three primary interactions:
@@
  • UE-to-AP Communication (Access Connectivity): UEs attach to a nearby AP to obtain Internet connectivity. This step uses standard link-layer and access-network protocols, such as IEEE 802.11 or 3GPP RRC/NAS for LTE/5G [40]. APs forward UE traffic to the Internet and do not participate in the decision-making process [30].
@@
  • UE-to-PSD Communication (Spectrum Querying): UEs submit spectrum-availability queries directly to PSDs using HTTPS over TLS to ensure confidentiality and integrity of location and device parameters. The request and response formats follow the JSON-based geolocation DB access protocol standardized by the IETF PAWS specification (RFC-7545) [30]. A typical UE-to-PSD exchange includes (i) a registration or device-descriptor message, (ii) a location-based AVAIL_CHANNELS query, and (iii) a PSD response containing permissible channels, power limits, and validity intervals. PSDs process each request using incumbent datasets and FCC rules before sending their response [40].
@@ 2.2 Cryptographic Building Blocks
$\bullet$ Delegatable Anonymous Credentials (DAC{DAC}): We use an attribute-based DAC scheme [32] that supports multi-level attribute issuance, controlled delegation, and anonymity. The main algorithms are summarized below (see [32]).
@@ A Delegatable Anonymous Credential (DAC) scheme is defined by the following algorithms:
$\bullet$ Distance Bounding Protocol (DBP): A DBP is a two-party protocol that verifies the physical proximity of two entities by measuring round-trip delays in a fast challenge–response exchange. We employ a PK-DBP [28] based on a one-pass authenticated key-agreement protocol using nonce Diffie–Hellman to derive a session key between the prover and verifier. This key is then used within the OTDB symmetric-DBP [43].
@@
$\bullet$ Revocable-iff-linked Linkable Ring Signature (RLRS): Ring signatures allow a user to anonymously sign on behalf of a group. An RLRS augments this by embedding an event identifier into each signature, enabling detection of multiple signatures produced by the same signer within that context and supporting controlled anonymity revocation when necessary [6].
@@ A Revocable-iff-linked Linkable Ring Signature (RLRS) scheme is a 6-tuple algorithm as shown below:
$\bullet$ Verifiable Delay Function (VDF): A VDF [8] enforces a prescribed delay by requiring sequential computation, while enabling a fast verification of the result. We use an RSA-based VDF [46], which provides compact proofs and constant-time verification.
@@ 3.1 Threat Model and Scope
  • Location Spoofing Attacks: Malicious users may attempt to bypass access restrictions by reporting incorrect location information to obtain spectrum resources or services beyond their authorized area. Such misuse can also arise from compromised or spoofed devices. Relevant threats include falsified coordinates, replayed location proofs, collusion with adversarial entities, distance fraud, relay or mafia fraud, and distance hijacking [34].
@@
Scope of Our Solution: The SLAPX framework is designed to preserve user anonymity and location privacy during spectrum access. It protects SUs while they query spectrum availability, obtain cryptographic puzzles, and access network services, mitigating threats such as identity exposure, location spoofing, and DoS attacks. Our work concentrates exclusively on the spectrum query and access phases. Privacy risks arising during user registration or from spectrum usage patterns (e.g., utilization inference) fall outside the scope of this study. Although the framework is tailored for SUs, it can be extended to support PUs interacting with PSD servers; however, PU-related protections are not considered here. Moreover, SLAPX does not address location leakage during ongoing spectrum use, user mobility or handover processes [21], nor does it cover timing-based attacks, side-channel leakage, or signal-based localization methods such as triangulation [7].
@@ 4 SLAPX: Design and Instantiation
Improvements over SLAP Framework and the SOTA: SLAPX extends our SLAP framework [15], incorporating substantial new material and the following key enhancements and additional features: (i) Comprehensive DoS countermeasures: We significantly strengthen DoS resilience by introducing on-demand, device-specific, and non-parallelizable VDFs, authenticated using lightweight standard signatures (e.g., ECDSA). This defense is further reinforced by a rate-limiting mechanism based on the linkability of event-oriented RLRSs, enabling effective DoS mitigation for both geolocation databases and service servers. (ii) Enhanced Proof of Location: We enhance the location proof mechanism by leveraging RLRS-based event linkability to enforce rate limiting on proof issuance and enable revocation of APs that collude with malicious users. This provides accountability and supports more realistic adversarial network settings. (iii) Comprehensive performance evaluation and network simulations: We substantially expand the evaluation by including detailed cryptographic benchmarking, comparisons with SOTA, and extensive network simulations that model real-world DoS and location spoofing scenarios under practical deployment conditions.
@@ 4.1 SLAPX Framework Initial Setup
Access Point Setup: Access Points (APs) within a geographical region jointly form a ring ($\mathcal{L}_{\texttt{ID}}$) that contains the ID of all participating APs. Each AP is provisioned with a cryptographic key pair $(\texttt{sk}_{\texttt{AP}},\texttt{PK}_{\texttt{AP}})$, generated by the FCC using the $\texttt{RLRS}{}.\texttt{Extract}{}(pp,\texttt{ID}_{\texttt{AP}})$ algorithm (Definition 3). APs periodically broadcast time-synchronized beacons $\beta_{\texttt{TS}}$ within predefined time windows TS to enable nearby device discovery. To verify UE proximity, each AP performs signal-based ranging using physical-layer measurements, including received signal strength (RSS) and round-trip time (RTT[3]. These measurements are processed via $\Delta\leftarrow\texttt{ProxVerify}{}(\texttt{RSS},\texttt{RTT},env_{\text{params}})$, which incorporates environment-specific factors such as path-loss models, multipath effects, and channel conditions to estimate the UE–AP distance and associated confidence bounds [31]. The resulting estimate $\Delta$ determines whether the UE lies within the AP’s trusted coverage region, providing a confidence interval that underpins subsequent authenticated access decisions. Additional details are provided in [38], with simulation results reported in Section 5.
@@ 5.1.1 Configuration and Experimental Setup
Configuration and Parameter Selection: We exclude signal transmission and network communication delays from our analysis, as these are typically constant-time operations and occur at the microsecond scale. Our configuration employs SHA-256 for hashing, structure-preserving set commitments [20], SPSEQ-UC signatures [32], and the BN-256 elliptic curve (EC) for binding and EC operations (with n=2048{n=2048}). Anonymous authentication is realized via Schnorr-type zero-knowledge proofs augmented with Damgård’s technique, following the DAC construction [32]. Non-interactive proofs are instantiated via the Fiat–Shamir transform, yielding an effective security level of approximately $100$ bits. The SGN digital signature is instantiated using ECDSA, a widely adopted and computationally efficient standard signature scheme.

### v2

@@ 1 Introduction
The rapid growth of wireless technologies, driven by mobile devices, Internet-of-Things (IoT) deployments, and emerging spectrum-hungry applications, has intensified pressure on traditionally static spectrum allocation policies [agarwal2022survey, chakraborty2023capow, 9321158]. To address spectrum underutilization, database-driven cognitive radio networks (DB-CRNs) and dynamic spectrum sharing have been introduced, enabling Secondary Users (SUs) to opportunistically access licensed spectrum owned by Primary Users (PUs) through regulatory-approved geolocation databases (DBs). A prominent example is the U.S. Citizens Broadband Radio Service (CBRS), which operates in the $3.5$ GHz band originally allocated to federal and satellite services [agarwal2022survey, grissa2021anonymous]. Moreover, dynamic spectrum sharing is expected to be a foundational element in future networks such as 6G systems [PATIL2024110697].
While this paradigm significantly improves spectrum efficiency, it also introduces a wide range of security and privacy challenges. In particular, regulatory authorities (e.g., FCC and ITU) require users to continuously report sensitive operational information, including precise location, device identity, and transmission parameters, to centralized DBs. This requirement raises serious concerns regarding user privacy, anonymity, and long-term profiling [jasim2021cognitive, grissa2016efficient]. Moreover, the location-centric design of DB-CRNs makes them inherently vulnerable to location spoofing attacks, in which malicious or compromised users falsify their positions to obtain unauthorized spectrum access [xin2016privacy]. At the same time, the centralized reliance on spectrum databases, the broadcast nature of wireless communication, and the proliferation of low-cost devices further expose DB-CRNs to denial-of-service (DoS) attacks that can disrupt spectrum coordination and degrade network availability [darzi2024privacy].
@@
    We develop and evaluate a novel and efficient framework, namely SLAPX, that leverages advanced cryptographic primitives to simultaneously address multiple, often conflicting, objectives in DB-CRNs. The proposed design achieves user location privacy and anonymity during spectrum querying and access while complying with enforced regulations. This paper extends our SLAP framework [^1] The preliminary version of our approach, namely SLAP, was presented and published at the IEEE SVCC 2025 Conference [darzi2025slap]., incorporating substantial new material and the following key enhancements and additional features. Designed for realistic adversarial settings, SLAPX remains robust against location spoofing and DoS attacks by malicious or compromised users. The key properties of SLAPX are outlined as follows: (i) Efficient Location Privacy-Preserving and Anonymous Spectrum Query: SLAPX provides strong anonymity and location privacy throughout the spectrum access workflow while remaining fully compliant with protocol and regulatory requirements. (ii) Adaptive Location Proof and Spoofing Resistance for DB-CRN: SLAPX comprises an adaptive location proof algorithm with dual scenario support and architectural flexibility that operates under realistic network assumptions with malicious or compromised users. (iii) DoS Mitigation with DB-CRN Architecture Compliance: SLAPX provides comprehensive DoS mitigation through a proactive defense mechanism that combines Verifiable Delay Functions (VDFs) with rate limiting enforced via the linkability properties of Revocable-iff-linked Linkable Ring Signatures (RLRSs).
@@
$\bullet$ Improvements over SLAP Framework: This paper extends our SLAP framework [^2] This paper is an extended version of our paper published at IEEE SVCC 2025 Conference [darzi2025slap], incorporating substantial new material and the following key enhancements and additional features:
@@ 2.1 System Architecture: DB-CRN
Network Components: The DB-CRN is a spectrum-sharing architecture in which unlicensed devices query trusted geolocation databases to obtain channel availability at their current location [PATIL2024110697]. This model was deployed in real FCC TV White Space (TVWS) systems and adopted in the Microsoft and Google database trials [caleffi2014database, das2015rfc]. The DB-CRN architecture consists of five key entities [order2023federal]:
@@
  • Spectrum Regulator: It is the regulatory authority that defines technical rules for DB-CRN operation, like incumbent protection requirements, device reporting and registration procedures, and transmit parameters [agarwal2022survey]. This entity authorizes and oversees geolocation database administrators and provides authoritative incumbent datasets and regulatory constraints. Throughout this manuscript, the FCC is used interchangeably to denote this spectrum regulatory authority for convenience.
@@
  • Servers: They represent external network or cloud services (e.g., CRN, cloud, edge nodes) that users access after obtaining spectrum authorization [grissa2016efficient]. They do not participate in spectrum assignment; instead, they simply carry user traffic over the channels approved by the PSD.
@@
  • Private Spectrum Databases (PSDs): PSDs are FCC-certified, cloud-based geolocation databases (e.g., Google, Microsoft, Spectrum Bridge) responsible for maintaining incumbent information, propagation constraints, and channel availability maps [chen2015protocol]. PSDs process spectrum-availability queries submitted by users and return permissible channels and power limits, synchronizing regularly according to FCC mandates [grissa2021anonymous].
@@
  • User Equipments (UEs): UEs are unlicensed devices of users (e.g., smartphones, sensors, and vehicular radios) that seek opportunistic access to licensed spectrum bands [grissa2016efficient]. Acting as SUs, they determine their location, submit channel-availability queries to PSDs, and operate strictly within the constraints specified in PSD responses to avoid interfering with PUs. Throughout this paper, the terms client and UE are used interchangeably.
@@
  • Access Points (APs): APs consist of WiFi access points or cellular base stations (LTE/5G macro or small cells) that provide the IP connectivity required for UEs to reach PSDs and servers over the Internet [darzi2025slap]. APs do not manage spectrum availability but offer backhaul connectivity and, in some advanced architectures, assist with localization or timing support.
@@
Communication Flow and Protocol Stack: In a DB-CRN, UEs communicate with PSDs over standard Internet paths to obtain spectrum availability, while servers and APs provide network connectivity and application-layer services. Communication exchanges follow an IP-based client–server model and use secure, lightweight protocols consistent with real FCC TVWS deployments and the IETF PAWS specification [liang2011cognitive]. The communication workflow consists of three primary interactions:
@@
  • UE-to-AP Communication (Access Connectivity): UEs attach to a nearby AP to obtain Internet connectivity. This step uses standard link-layer and access-network protocols, such as IEEE 802.11 or 3GPP RRC/NAS for LTE/5G [thippeswamy2016physical]. APs forward UE traffic to the Internet and do not participate in the decision-making process [liang2011cognitive].
@@
  • UE-to-PSD Communication (Spectrum Querying): UEs submit spectrum-availability queries directly to PSDs using HTTPS over TLS to ensure confidentiality and integrity of location and device parameters. The request and response formats follow the JSON-based geolocation DB access protocol standardized by the IETF PAWS specification (RFC-7545) [liang2011cognitive]. A typical UE-to-PSD exchange includes (i) a registration or device-descriptor message, (ii) a location-based AVAIL_CHANNELS query, and (iii) a PSD response containing permissible channels, power limits, and validity intervals. PSDs process each request using incumbent datasets and FCC rules before sending their response [thippeswamy2016physical].
@@ 2.2 Cryptographic Building Blocks
$\bullet$ Delegatable Anonymous Credentials (DAC{DAC}): We use an attribute-based DAC scheme [mir2023practical] that supports multi-level attribute issuance, controlled delegation, and anonymity. The main algorithms are summarized below (see [mir2023practical]).
@@ A Delegatable Anonymous Credential (DAC) scheme is defined by the following algorithms:
$\bullet$ Distance Bounding Protocol (DBP): A DBP is a two-party protocol that verifies the physical proximity of two entities by measuring round-trip delays in a fast challenge–response exchange. We employ a PK-DBP [kilincc2016efficient] based on a one-pass authenticated key-agreement protocol using nonce Diffie–Hellman to derive a session key between the prover and verifier. This key is then used within the OTDB symmetric-DBP [vaudenay2015private].
@@
$\bullet$ Revocable-iff-linked Linkable Ring Signature (RLRS): Ring signatures allow a user to anonymously sign on behalf of a group. An RLRS augments this by embedding an event identifier into each signature, enabling detection of multiple signatures produced by the same signer within that context and supporting controlled anonymity revocation when necessary [au2013secure].
@@ A Revocable-iff-linked Linkable Ring Signature (RLRS) scheme is a 6-tuple algorithm as shown below:
$\bullet$ Verifiable Delay Function (VDF): A VDF [boneh2018verifiable] enforces a prescribed delay by requiring sequential computation, while enabling a fast verification of the result. We use an RSA-based VDF [wesolowski2020efficient], which provides compact proofs and constant-time verification.
@@ 3.1 Threat Model and Scope
  • Location Spoofing Attacks: Malicious users may attempt to bypass access restrictions by reporting incorrect location information to obtain spectrum resources or services beyond their authorized area. Such misuse can also arise from compromised or spoofed devices. Relevant threats include falsified coordinates, replayed location proofs, collusion with adversarial entities, distance fraud, relay or mafia fraud, and distance hijacking [nguyen2019spoofing].
@@
Scope of Our Solution: The SLAPX framework is designed to preserve user anonymity and location privacy during spectrum access. It protects SUs while they query spectrum availability, obtain cryptographic puzzles, and access network services, mitigating threats such as identity exposure, location spoofing, and DoS attacks. Our work concentrates exclusively on the spectrum query and access phases. Privacy risks arising during user registration or from spectrum usage patterns (e.g., utilization inference) fall outside the scope of this study. Although the framework is tailored for SUs, it can be extended to support PUs interacting with PSD servers; however, PU-related protections are not considered here. Moreover, SLAPX does not address location leakage during ongoing spectrum use, user mobility or handover processes [gao2012location], nor does it cover timing-based attacks, side-channel leakage, or signal-based localization methods such as triangulation [bahrak2014protecting].
@@ 4 SLAPX: Design and Instantiation
Improvements over SLAP Framework and the SOTA: SLAPX extends our SLAP framework [darzi2025slap], incorporating substantial new material and the following key enhancements and additional features: (i) Comprehensive DoS countermeasures: We significantly strengthen DoS resilience by introducing on-demand, device-specific, and non-parallelizable VDFs, authenticated using lightweight standard signatures (e.g., ECDSA). This defense is further reinforced by a rate-limiting mechanism based on the linkability of event-oriented RLRSs, enabling effective DoS mitigation for both geolocation databases and service servers. (ii) Enhanced Proof of Location: We enhance the location proof mechanism by leveraging RLRS-based event linkability to enforce rate limiting on proof issuance and enable revocation of APs that collude with malicious users. This provides accountability and supports more realistic adversarial network settings. (iii) Comprehensive performance evaluation and network simulations: We substantially expand the evaluation by including detailed cryptographic benchmarking, comparisons with SOTA, and extensive network simulations that model real-world DoS and location spoofing scenarios under practical deployment conditions.
@@ 4.1 SLAPX Framework Initial Setup
Access Point Setup: Access Points (APs) within a geographical region jointly form a ring ($\mathcal{L}_{\texttt{ID}}$) that contains the ID of all participating APs. Each AP is provisioned with a cryptographic key pair $(\texttt{sk}_{\texttt{AP}},\texttt{PK}_{\texttt{AP}})$, generated by the FCC using the $\texttt{RLRS}{}.\texttt{Extract}{}(pp,\texttt{ID}_{\texttt{AP}})$ algorithm (Definition 3). APs periodically broadcast time-synchronized beacons $\beta_{\texttt{TS}}$ within predefined time windows TS to enable nearby device discovery. To verify UE proximity, each AP performs signal-based ranging using physical-layer measurements, including received signal strength (RSS) and round-trip time (RTT[abuyaghi2024positioning]. These measurements are processed via $\Delta\leftarrow\texttt{ProxVerify}{}(\texttt{RSS},\texttt{RTT},env_{\text{params}})$, which incorporates environment-specific factors such as path-loss models, multipath effects, and channel conditions to estimate the UE–AP distance and associated confidence bounds [lopez2019ieee]. The resulting estimate $\Delta$ determines whether the UE lies within the AP’s trusted coverage region, providing a confidence interval that underpins subsequent authenticated access decisions. Additional details are provided in [rosler2025improving], with simulation results reported in Section 5.
@@ 5.1.1 Configuration and Experimental Setup
Configuration and Parameter Selection: We exclude signal transmission and network communication delays from our analysis, as these are typically constant-time operations and occur at the microsecond scale. Our configuration employs SHA-256 for hashing, structure-preserving set commitments [fuchsbauer2019structure], SPSEQ-UC signatures [mir2023practical], and the BN-256 elliptic curve (EC) for binding and EC operations (with n=2048{n=2048}). Anonymous authentication is realized via Schnorr-type zero-knowledge proofs augmented with Damgård’s technique, following the DAC construction [mir2023practical]. Non-interactive proofs are instantiated via the Fiat–Shamir transform, yielding an effective security level of approximately $100$ bits. The SGN digital signature is instantiated using ECDSA, a widely adopted and computationally efficient standard signature scheme.